OpenStudy (anonymous):

Solve the equation using the zero-product property. (x-2)(x+7)=0 A. 2,-7 B. 2,7 C. -1,1 D. -2,-7

OpenStudy (anonymous):

either x-2 =0 OR x+7 = 0

OpenStudy (anonymous):

x=2 , x=-7 these make the inside 0

OpenStudy (anonymous):

of the parentheses

Parth (parthkohli):

Zero-product property says:\[(\color{Red}{number})(\color{blue}{number}) = 0 \implies\color{blue}{number}=0 \quad or \quad\color{red}{number }=0 \]

OpenStudy (anonymous):

now solve for x for both the equation and get the value of x

OpenStudy (anonymous):

Thanks guys :)

Parth (parthkohli):

Therefore:\[ (x - 2)(x + 7) = 0 \implies x - 2= 0\quad \textbf{or} \quad x + 7 =0\]

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

thanks